OpenStudy (anonymous):

please help ill do anything Write an equation in slope-intercept form of the line that passes through the points (-1,-4) and (1,4)? write the equation in slope-intercept form

OpenStudy (anonymous):

first you need to be able to find the slope between two points then take that slope and chose any of the two points you are given and plug them into the slope intercept form.

OpenStudy (anonymous):

that's where im confused I have the slope 8/-2

OpenStudy (anonymous):

4-(-4) / (1-(-1) = 8/2 = 4

OpenStudy (anonymous):

what you wiling to do for the answer? ;)... this involves using the equation: Y-y = m(X-x) where you have to use one of the points as the y and the x of the equation.

OpenStudy (anonymous):

\[y-(4) =4(x-1)\]

OpenStudy (anonymous):

\[y= 4x+0\]